# Efficiently Vectorized MCMC on Modern Accelerators

### v1

@@ 4.1 The long-run relative cost of the FSM
Substituting the form of the constants into Theorem 4.1, the ratio of the long-run expected runtimes, $E(m)$, is given by the ratio
@@
and that this bound is tight. We refer to $R(m)$ as the ‘theoretical efficiency bound’ for the FSM. Note from Equation 18 that minimizing $\alpha$ and $K$ improves the efficiency $E(m)$ of the FSM. In Section 5 we will introduce two techniques to minimize $\alpha$ and $K$ in practice, which enable us to nearly obtain the efficiency bound $R(m)$ for certain MCMC algorithms.
The scale of potential speed-ups. The size of $R(m)$ depends (only) on the underlying distribution of $N_{1}$ (since $N_{i}=_{d}N_{j}$ $\forall i,j\in[m]$). Whenever $N_{1}$ is sub-exponential, it is known that $R(m)=\mathcal{O}(\ln(m))$ (Vershynin,, 2018). Although this implies a slow rate of increase in $m$, $R(m)$ can be still be very large for small values of $m$. For example, if $N_{j}/B\sim Bern(p)$ (i.e., one either needs zero or $B$ iterations to get a sample), then $R(m)=(1-(1-p)^{m})/p$ and converges to $1/p$ exponentially fast as $m$ increases. For small $p$ this can be very large even for small $m$. In general $R(m)$ is sensitive to the skewness of the distribution: distributions on $[0,B]$ with zero skewness have $R(m)=2$, whilst distributions with skew of 10 can have $R(m)\approx 100$; see Figure 4. Intuitively, these are the distributions where chains are slow only occasionally, but at least one chain is slow often. In such cases, our FSM-design can lead to enormous efficiency gains, as we show in experiments.
@@ 5.1 Step bundling to reduce $K$
Given an FSM with step function step defined by code blocks $S_{1},...,S_{K}$ and transition function $\delta$, one can ‘bundle’ multiple FSM steps together using a modified step function, bundled_step, which replaces the switch over the algorithm state $k$ in step with a series of separate conditional statements. This is shown in Algorithm 4 for an example with two state functions. Note that under the “run all branches and mask” behaviour of vmap, vmap(step) and vmap(bundled_step) have the same cost. However, whenever bundled_step runs $S_{1}$ and $\delta$ returns $k=2$, it immediately also runs $S_{2}$. This essentially reduces the ‘effective’ number of states $K$ and/or reduces the overall number of steps needed to recover a sample, increasing efficiency. In principle, the block ordering can be optimized for sequences that are expected to occur with higher probability.
@@
5.2 Cost amortization
@@
Note that now $g(\theta)$ is only called once per step when using vmap on amortized_step, even if it required for every code block $S_{k}$. This can extend to multiple functions $g_{1},..,g_{l}$. In practice, we find amortization most powerful for when the log-density is expensive and needed in multiple state functions, in which case we can set $g(\cdot)=\log p(\cdot)$. The elliptical slice sampler is one such case, where the log-pdf is needed in both $S_{1}$ (i.e. the block before the while loop - INIT in Figure 3) and $S_{2}$ (i.e. the block for the body of the while loop - SHRINK in Figure 3), to check the while loop terminator.
@@ 6   Related work
FSMs in Machine Learning. Previous work in machine learning has used the framework of FSMs to design image-based neural networks (Ardakani et al.,, 2020) and Bayesian non-parametric time series models (Ruiz et al.,, 2018), as well as extract representations from Recurrent-Neural-Networks (RNNs) (Muškardin et al.,, 2022; Cechin et al.,, 2003; Tiňo et al.,, 1998; Zeng et al.,, 1993; Koul et al.,, 2018; Svete and Cotterell,, 2023). To our knowledge, our work is the first that use used FSMs as a framework to represent MCMC algorithms, and design novel implementations.
Efficient MCMC on Modern Hardware. Given the recursive nature of HMC-NUTS, previous work has reformulated the algorithm for compatibility with machine learning frameworks that cannot naively support recursion (Abadi et al.,, 2016; Phan et al.,, 2019; Lao et al.,, 2020). However, these implementations do not address synchronization inefficiencies caused by automatic vectorization tools. Our work bears some similarities with a general-purpose algorithm proposed in the High-Performance-Computing literature for executing batched recursive programs (Radul et al.,, 2020). However whilst both their method and ours breaks programs down into smaller blocks for efficient vectorization, there are several major differences. Our approach provides a recipe for implementing a range of single-chain iterative MCMC algorithms and is fully compatible with automatic vectorization tools like vmap. In contrast, their algorithm is designed to work with recursive programs and requires code which is already batched. Crucially, in order to avoid synchronization barriers due to while loops, this code cannot have been batched with a vmap-equivalent vectorization tool, because vmap converts while loops into a single batched primitive which cannot be broken down by their algorithm. Additionally, our method breaks programs down into the coarsest (while-loop-free) blocks possible and allows for control flow within each block, which we showed is crucial for optimal performance in frameworks which run and mask branches. Their algorithm does not allow blocks to contain any control-flow, yielding very small blocks. This can be detrimental to performance under the default “run and mask” assumption made in their work. By focusing on MCMC, we also obtain provable speed-ups under appropriate statistical settings and FSM design.
@@ 7   Experiments
Figure 5: Mean and standard deviation walltimes (LHS) and ESS (RHS) of the symmetric Delayed-Rejection Algorithm (Mira et al.,, 2001) using Algorithm 1 and our FSM implementation Algorithm 3 (both with vmap), on a univariate Gaussian (10 random seeds). To test the effect of the number of states and step function bundling on performance, FSM uses step with the while loop body (PROPOSE in Figure 3) split into $4$ states, whilst FSM-condensed uses bundled_step with the same states. By bundling the steps together, we achieve a $\sim 3$x efficiency gain.
The following experiments evaluate FSM implementations of different MCMC algorithms with while loops against their standard (non-FSM) implementations, as well as other MCMC methods in one experiment. All methods (including ours) consist of single chain MCMC algorithms written in JAX, turned into multi-chain methods with vmap, and compiled using jit. All experiments are run in JAX on a NVIDIA A100 GPU with 32GB CPU memory.
@@ 7.1 Delayed-Rejection MH on a Simple Gaussian
We first illustrate basic properties of the FSM conversion on a toy example. The MCMC algorithm used is symmetric Delayed-Rejection Metropolis Hastings (DRMH) (Mira et al.,, 2001), which is a simple example of a delayed rejection method. Symmetric DRMH modifies the Random-Walk Metropolis-Hastings algorithm by iteratively re-centering the proposal distribution on the rejected sample and resampling until either acceptance occurs or a maximum number of tries $M$ is reached. To ensure detailed balance hods, the acceptance probability formula is adjusted at each step to account for past rejections.
Experimental setup. As a toy problem, we implement symmetric DRMH using (vmap’ed) Algorithm 1 (baseline) and Algorithm 3 (ours) to sample from a univariate Gaussian $\mathcal{N}(0,1)$, varying the number of chains. We use a $\mathcal{N}(x,0.1)$ proposal distribution with $M=100$ tries per sample and draw 10,000 samples per chain. Although DRMH has 3 state functions by default (see Figure 3), the INIT and DONE states are essentially empty, and so just a single state function can be used for the while loop body. This means an appropriate FSM implementation should be able to get close to $R(m)$, up to overheads. To illustrate the importance of designing FSMs with as few (effective) state functions as possible, we implement an FSM which unrolls the while loop body into four different state functions, as well as a (condensed) FSM which uses step function bundling to effectively use a single state.
@@ 7.2 Elliptical Slice Sampling on Real Estate Data
Figure 7: Left: contours of the first two dimensions of a correlated mixture of Gaussians, along with a single chain of HMC-NUTS and MALA. Middle Left: Histogram of the number of integration steps taken per sample for a single NUTS chain. Middle Right: histogram of the maximum number of integration steps taken per sample across 500 chains. Right: Effective samples per minute for the standard BlackJAX HMC-NUTS implementation, and our FSM implementation of HMC-NUTS (average and standard error bars from 5 seeds). The FSM achieves speed ups of nearly an order of magnitude for 100 chains, and or than half an order of magnitude for 500 chains.
@@
7.3 HMC-NUTS on a high-dimensional correlated MoG
@@
Experimental setup. We implement NUTS on a 100-dimensional correlated mixture of Gaussians ($\rho=0.99$), with the mixture modes placed along the principal direction at $(-5\cdot\bm{1},\bm{0},5\cdot\bm{1})$. We use a pre-tuned step-size with acceptance rate $\sim 0.85$ and set $M=I$ for the mass matrix. We draw $n=1000$ samples per chain and vary ##\## chains $m$.
@@
Results. The contours of the log-density and a trajectory of a single NUTS chain (1000 samples) are displayed in Figure 7 (one dot = one sample). The typical distance traveled by NUTS is small (few integration steps), with the occasional large jump (many integration steps) when the momentum sample aligns with the principal direction. The FSM yields speedups of nearly an order of magnitude for $m=100$, and about half an order of magnitude for $m=500$. This is reflects the fact that the marginal distribution of integration steps is very skewed (i.e. $R(m)$ is large). As Figure 7 shows, the probability a sample needs less than $20$ steps is $\sim 0.95$ and needs $>1000$ steps is $\sim 0.01$. However, the probability that at least one chain needs more than 1000 steps is $\sim 0.99$. Note that one can avoid sychronization barriers and obtain very high ESS/Sec using a simpler algorithm like MALA, but this fails to explore the distribution (LHS Figure 7).
@@
7.4 Transport Elliptical Slice Sampling for distributions with challenging local geometry
Table 1: ESS/Second for different methods on the four benchmark problems used in Cabezas and Nemeth, (2023).
MEADS CHEES NeuTra TESS TESS-FSM
Predator Prey 1.53 nan 1.59 2.27 3.80
Google Stock 480.76 60.19 185.12 1116.31 2426.16
German Credit 141.50 198.99 186.17 58.95 59.04
BOD 64.846 247.02 130.59 2978.03 3252.64
@@ 7.4 Transport Elliptical Slice Sampling for distributions with challenging local geometry
Transport elliptical slice sampling (TESS), due to Cabezas and Nemeth, (2023), is a variant of elliptical slice sampling designed for certain challenging local geometries. It essentially uses a normalizing flow $T$ to ‘precondition’ the distribution $\pi$, and does elliptical slice sampling on the transformed distribution $T_{\#}\pi$. Since $T$ is learned to approximately map $\pi$ to $\mathcal{N}(0,I)$, the geometry of $T_{\#}\pi$ makes it much easier to sample from. Once these samples are obtained, they can be pushed through $T^{-1}$ to recover samples from $\pi$. TESS achieves particularly good results on distributions with ‘funnel’ geometries, with which gradient-based methods like HMC tend to struggle (Gorinova et al.,, 2020).
@@
Experimental setup. TESS has similar overall structure as the elliptical slice sampler, and conversion results in the same ‘single loop’ FSM in Figure 3. We compare TESS with and without FSM conversion on the four benchmark sampling problems used in Cabezas and Nemeth, (2023), which are chosen for their challenging geometries. As baselines we use two recently proposed adaptive HMC variants (MEADS (Hoffman and Sountsov,, 2022), and CHEES (Hoffman et al.,, 2021)) as well as NeuTra (Hoffman et al.,, 2019), which uses a similar preconditioning strategy to TESS, but with HMC as the sampling algorithm. For all methods, we run 128 chains of 1000 samples each, with each algorithm’s hyperparameters pre-tuned using 400 warm-up steps.
Results. In all four cases, TESS-FSM improved ESS/second over TESS. In three these cases TESS-FSM achieved the best ESS/second, and in two cases the speed-up over TESS was roughly by an order of 2. Since all methods considered here are state-of-the-art (SOTA) approaches to sampling from distributions with challenging geometries, this demonstrates our FSM implementation method can be used to improve SOTA performance on such tasks.
@@ Appendix A Mathematical appendix
Let $(Z_{i})_{i\geq 1}$ be a Markov Chain with measurable state space $\mathcal{Z}$ and stationary distribution $\pi$, and let $f:\mathcal{Z}\to[0,B]$ be measurable and bounded. Then, for any $\epsilon>0$,
@@
where $\lambda$ is the spectral gap of $\pi$ (as defined in Fan et al., (2021)) which quantifies the speed of convergence of the Markov chain towards $\pi$ (Rudolf,, 2011).
@@
Appendix B Additional Details and Results

### v2

@@ 4.2 The Relative Long-Run Cost of the FSM
Using the form of the constants in Theorem 4.1, the ratio of the long-run costs, $E(m):=C_{0}(m)/C_{F}(m)$, is given by
@@
and that this bound is tight. We refer to $R(m)$ as the ‘theoretical efficiency bound’ for the FSM. Note from Equation 16 that minimizing $\alpha$ and $K$ improves the efficiency $E(m)$ of the FSM. In Section 5 we introduce two techniques to minimize $\alpha$ and $K$ in practice, which enable us to nearly obtain the efficiency bound $R(m)$ for certain MCMC algorithms.
Scale of Potential Speed-ups. The size of $R(m)$ depends (only) on the underlying distribution of $N_{1}$ (since $N_{i}=_{d}N_{j}$ $\forall i,j\in[m]$). Whenever $N_{1}$ is sub-exponential, it is known that $R(m)=\mathcal{O}(\ln(m))$ (Vershynin,, 2018). Although this implies a slow rate of increase in $m$, $R(m)$ can be still be very large for small values of $m$. For example, if $N_{j}/B\sim Bern(p)$ (i.e., one either needs zero or $B$ iterations to get a sample), then $R(m)=(1-(1-p)^{m})/p$ and converges to $1/p$ exponentially fast as $m$ increases. For small $p$ this can be very large even for small $m$. In general $R(m)$ is sensitive to the skewness of the distribution: distributions on $[0,B]$ with zero skewness have $R(m)=2$, whilst distributions with skew of 10 can have $R(m)\approx 100$; see Figure 4. Intuitively, these are the distributions where chains are slow only occasionally, but at least one chain is slow often. In such cases, our FSM-design can lead to enormous efficiency gains, as we show in experiments.
@@ 5.1 Step Bundling to Reduce $K$
Given an FSM with step defined by code blocks $S_{1},...,S_{K}$ and transition function $\delta$, one can ‘bundle’ multiple steps together using a modified step function, bundled_step, which replaces the switch over the algorithm state $k$ in step with a series of separate conditional statements. This is shown in Algorithm 4 for an example with two state functions. Note that under the “run all branches and mask” behaviour of vmap, vmap(step) and vmap(bundled_step) have the same cost. However, whenever bundled_step runs $S_{1}$ and $\delta$ returns $k=2$, it immediately also runs $S_{2}$. This reduces the ‘effective’ number of states $K$ and/or the overall number of steps needed to recover a sample, increasing efficiency. In principle, the block ordering can be optimized for sequences that are expected to occur with higher probability. However, we found chronological order a surprisingly effective heuristic.
@@
5.2 Cost Amortization to Reduce $\alpha$
@@
Note the function $g$ is now only called once per step when using vmap on amortized_step. In particular, if the executions of $g$ cost (in total) $\beta\in[0,1]$ fraction of the total cost of all blocks, amortization will reduce this fraction to $\frac{\beta}{M}+(1-\beta)$ We indeed observe this when amortizing $g=\log p$ for the elliptical slice sampler in Section 7.1. In that scenario, $M=2$ and $\beta\approx 1$ as the log-likelihood is very expensive, resulting in a 2×2$\times$2 × speed-up over no amortization. One issue is that amortization only allows steps to be bundled which do not require calls of $g$. We adapt the algorithm to handle this in Appendix B.
@@ 6 Related work
FSMs in Machine Learning. Previous work in machine learning has used the framework of FSMs to design image-based neural networks (Ardakani et al.,, 2020) and Bayesian nonparametric time series models (Ruiz et al.,, 2018), as well as extract representations from Recurrent-Neural-Networks (RNNs) (Muškardin et al.,, 2022; Cechin et al.,, 2003; Tiňo et al.,, 1998; Zeng et al.,, 1993; Koul et al.,, 2018; Svete and Cotterell,, 2023). To our knowledge, our work is the first that uses FSMs as a framework to represent MCMC algorithms, and design novel implementations.
Efficient MCMC on Modern Hardware. Given the recursive nature of HMC-NUTS, previous work has reformulated the algorithm for compatibility with machine learning frameworks that cannot naively support recursion (Abadi et al.,, 2016; Phan et al.,, 2019; Lao et al.,, 2020). However, these implementations do not address synchronization inefficiencies caused by automatic vectorization tools. Our work bears some similarities with a general-purpose algorithm proposed in the High-Performance-Computing literature for executing batched recursive programs (Radul et al.,, 2020). Both their method and ours breaks programs down into smaller blocks for efficient vectorization, but there are several major differences. Our approach provides a recipe for implementing single-chain iterative MCMC algorithms and is fully compatible with automatic vectorization tools like vmap. In contrast, their algorithm is designed for recursive programs and requires code which is already batched. Crucially, to avoid synchronization barriers due to while loops, this code cannot have been batched with a vmap-equivalent vectorization tool, since vmap converts while loops into a single batched primitive that cannot be broken down by their algorithm. Additionally, our method uses the fewest (while-loop-free) blocks possible and allows for control flow within blocks (which we showed is crucial for optimal performance under the “run and mask” paradigm). By contrast, their algorithm does not allow blocks to contain any control-flow, yielding many small blocks. We also obtain provable speed-ups for MCMC via theoretical analysis.
@@ 7 Experiments
Figure 6: Mean and standard deviation walltimes (LHS) and ESS (RHS) of the symmetric Delayed-Rejection Algorithm (Mira et al.,, 2001) using Algorithm 1 and our FSM implementation Algorithm 3 (both with vmap), on a univariate Gaussian (10 random seeds). FSM uses step with the while loop body (PROPOSE in Figure 3) split into $4$ states, whilst FSM-condensed uses a single step. Using a single step leads to a $\sim 3$x efficiency gain.
The following experiments evaluate FSMs for different MCMC algorithms with while loops against their standard (non-FSM) implementations, as well as other MCMC methods in one experiment. All methods consist of single chain MCMC algorithms written in JAX, turned into multi-chain methods with vmap, and compiled using jit. All experiments are run in JAX on an NVIDIA A100 GPU with 32GB CPU memory. Code can be found at https://github.com/hwdance/jax-fsm-mcmc.
@@ 7.1 Delayed-Rejection MH on a Simple Gaussian
We first illustrate basic properties of the FSM conversion on a toy example. The MCMC algorithm used is symmetric Delayed-Rejection Metropolis Hastings (DRMH) (Mira et al.,, 2001), which is a simple example of a delayed rejection method. Symmetric DRMH modifies the Random-Walk Metropolis-Hastings algorithm by iteratively re-centering the proposal distribution on the rejected sample and resampling until either acceptance occurs or a maximum number of tries $M$ is reached. To ensure detailed balance, the acceptance probability is adjusted to account for past rejections.
Experimental setup. We implement symmetric DRMH using (vmap’ed) Algorithm 1 (baseline) and Algorithm 3 (ours) to sample from a univariate Gaussian $\mathcal{N}(0,1)$, varying the number of chains. We use a $\mathcal{N}(x,0.1)$ proposal distribution with $M=100$ tries per sample and draw 10,000 samples per chain. Although DRMH has 3 state functions by default (see Figure 3), the INIT and DONE states are essentially empty, and so just a single state function can be used for the while loop body. This means an appropriate FSM implementation should be able to get close to $R(m)$, up to overheads. To illustrate the importance of designing FSMs with as few (effective) state functions as possible, we implement an FSM which unrolls the while loop body into four different state functions, as well as a (condensed) FSM with a single state function (i.e. effectively uses bundling).
@@ 7.2 Elliptical Slice Sampling on Real Estate Data
Figure 7: Left: contours of the first two dimensions of a correlated mixture of Gaussians, along with a single chain of HMC-NUTS and MALA. Middle Left: Histogram of the number of integration steps taken per sample for a single NUTS chain. Middle Right: histogram of the maximum number of integration steps taken per sample across 500 chains. Right: Effective samples per minute for the standard BlackJAX HMC-NUTS implementation, and our FSM implementation of HMC-NUTS (average and standard error bars from 5 seeds). The FSM achieves speed ups of nearly an order of magnitude for 100 chains, and more than half an order of magnitude for 500 chains.
@@
7.3 HMC-NUTS on High-Dimensional Correlated MoN
@@
Experimental setup. We implement NUTS on a 100-dimensional correlated mixture of Gaussians ($\rho=0.99$), with the mixture modes placed along the principal direction at $(-10\cdot\bm{1},\bm{0},10\cdot\bm{1})$. We use a pre-tuned step-size with acceptance rate $\sim 0.85$ and identity mass matrix. We draw $n=1000$ samples per chain and vary ##\## chains $m$.
@@
Results. The trajectory of a single NUTS chain (1000 samples) are displayed on the LHS of Figure 7 (one dot = one sample). The typical distance traveled by NUTS is small (few integration steps), with the occasional large jump (many integration steps) when the momentum sample aligns with the principal direction. In particular, the probability a sample requires less than $20$ integration steps is $\sim 0.95$ and needs $>1000$ steps is $\sim 0.01$, but the probability that at least one chain needs more than 1000 steps is $\sim 0.99$ (Middle Figure 7). This results in FSM speed-ups of nearly an order of magnitude for $m=100$, and about half an order of magnitude for $m=500$ (RHS Figure 7). Note that one can avoid sychronization barriers and obtain very high ESS/Sec using a simpler algorithm like MALA, but this fails to explore the distribution (LHS Figure 7).
@@
7.4 Transport Elliptical Slice Sampling and NUTS on Distributions with Challenging Local Geometry
Table 1: Left: ESS/sec for baselines on Predator Prey (PP), Google Stock (GS), German Credit (GC), Biochemical Oxygen Demand (BOD) distributions. Right: FSM speed-ups for NUTS/TESS.
ESS/sec (non‐FSM) FSM Speed‐up
Dist. MEADS CHEES NUTS TESS NUTS TESS
PP 1.5 nan 0.02 2.3 1.5× 1.7×
GS 480.8 60.2 0.15 1116 3.5× 2.2×
GC 141.5 199.0 1.71 58.95 1.2× 1.0×
BOD 64.85 247.0 0.56 2978 0.8× 1.1×
@@ 7.4 Transport Elliptical Slice Sampling and NUTS on Distributions with Challenging Local Geometry
Transport elliptical slice sampling (TESS), due to Cabezas and Nemeth, (2023), is a state-of-the-art variant of elliptical slice sampling designed for challenging local geometries. It uses a normalizing flow $T$ to ‘precondition’ the distribution $\pi$ into an approximate Gaussian, does elliptical slice sampling on the transformed distribution $T_{\#}\pi$, and then pushes generated samples through $T^{-1}$ to recover samples from $\pi$. TESS achieves particularly good results on distributions with ‘funnel’ geometries, on which gradient-based methods like HMC-NUTS tend to struggle (Gorinova et al.,, 2020). TESS is similar to the elliptical slice sampler and so the FSM has the same ‘single loop’ structure (see Figure 3).
@@
Experimental setup We examine the speed-ups using FSMs for TESS and NUTS on the four benchmark distributions in Cabezas and Nemeth, (2023), which are chosen for their challenging geometries. As baselines we use two state-of-the-art adaptive HMC variants (MEADS (Hoffman and Sountsov,, 2022), and CHEES (Hoffman et al.,, 2021)). We expect NUTS to perform poorly on these problems (as observed in Cabezas and Nemeth, (2023)), but implementing it lets us at least examine the FSM speed-ups. For all methods, we average results over 128 chains of 1000 samples, with hyperparameters pre-tuned over 400 warm-up steps.
Results are in Table 1. TESS-FSM improved ESS/sec over TESS in all cases (in 2/4 cases by 2$\sim$ 2$\times$ 2 ×) and in 3/4 cases improved the best ESS/sec. NUTS performed poorly as expected, but the FSM improved ESS/sec in 3/4 cases, and >3×>\!3$\times$> 3 × in one case. Speed-ups for NUTS and TESS were larger for tasks with an expensive log-pdf (PP, GS). This is because most time is spent in the loop states where the log-pdf is called, and a more costly log-pdf reduces the gap between the cost of executing these states for non-FSMs ($c_{k}$) and FSMs ($\sum_{j=1}^{m}c_{j}$) (see Section 4), improving efficiency.
Acknowledgements
@@
Impact Statement
This work improves the execution efficiency of MCMC algorithms. Our method has the potential to significantly reduce time, cost, and energy expenditure for a range of scientific applications.
@@ Appendix A Mathematical Appendix
We first define formally the notion of an absolute spectral gap, from (Fan et al.,, 2021).
Definition A.1 (Absolute Spectral Gap).
Let $\mathcal{Z}$ and $\pi$ denote the state space and the invariant measure of a Markov chain $\{Z_{i}\}_{i\geq 1}$. For a function $f:\mathcal{Z}\to\mathbb{R}$, write $\pi(h)\;:=\;\int h(z)\,\pi(dz)$. Let $L_{2}(\pi)\;=\;\{\,h:\pi(h^{2})<\infty\}$ be the Hilbert space consisting of $\pi$-square-integrable functions, and $L_{2}^{0}(\pi)\;=\;\{\,h\in L_{2}(\pi):\pi(h)=0\}$ be its subspace of $\pi$-meanzero functions. The transition probability kernel of the Markov chain, denoted by $P$, is viewed as an operator acting on $L_{2}(\pi)$. Let $\lambda\in[0,1]$ be the operator norm of $P$ acting on $L_{2}^{0}(\pi)$. Then, $1-\lambda$ is the absolute spectral gap of the Markov chain.
@@
Let $(Z_{i})_{i\geq 1}$ be a Markov Chain with measurable state space $\mathcal{Z}$, stationary distribution $\pi$, and absolute spectral gap $1-\lambda\in(0,1]$. Let $f:\mathcal{Z}\to[0,B]$ be measurable and bounded. Then, for any $\epsilon>0$,
@@
where $1-\lambda$ is the absolute spectral gap of $\pi$.
@@
Appendix B Additional Details